OpenStudy (goldrush18):

find the equation of the line that is perpendicular to the line y=3x+2, and passes through the point (0,1) find the equation of the line that is perpendicular to the line y=3x+2, and passes through the point (0,1) @Mathematics

OpenStudy (anonymous):

perpendicular means that the slope is negative reciprocal of the original slope new slope is -1/3 1 = -0/3 + b b = 1 y = -x/3 + 1

OpenStudy (goldrush18):

what does b stand for

OpenStudy (goldrush18):

kinda confused

OpenStudy (anonymous):

b is your y-intercept